OpenStudy (anonymous):

6. The force exerted on a spring is proportional to the distance the spring is stretched from its relaxed position. Suppose you stretch a spring a distance of d inches by applying a force of F pounds. For your spring, d/F = 0.8. You apply forces between 40 lb and 54 lb, inclusive. What inequality describes the distances the spring is stretched? (1 point)

OpenStudy (anonymous):

totally lost please help

OpenStudy (anonymous):

SInce d is proportional to F :: d = k*F For your spring \[\frac{ d }{ F } = \frac{ 8 }{10 }\] so we can set up the equation d = k*10 Therefore k = 0.8 for your spring. To get your distance interval since you have the Force interval and K; all you do now is solve for both end points. Remember d = k*F For F = 40 Ib d = 0.8 * 40 = 32 m For F = 54 Ib = = 0.8 * 54 = 43.2 m So your inequality = \[32 \le d \le 43.2\]

OpenStudy (anonymous):

Correction: So we set up the equation 8 = k*10 and solve for k

OpenStudy (anonymous):

thanks so much!